# Transformers as Unsupervised Learning Algorithms: A study on Gaussian Mixtures

### v1

@@ Abstract
The transformer architecture has demonstrated remarkable capabilities in modern artificial intelligence, among which the capability of implicitly learning an internal model during inference time is widely believed to play a key role in the under standing of pre-trained large language models. However, most recent works have been focusing on studying supervised learning topics such as in-context learning, leaving the field of unsupervised learning largely unexplored. This paper investigates the capabilities of transformers in solving Gaussian Mixture Models (GMMs), a fundamental unsupervised learning problem through the lens of statistical estimation. We propose a transformer-based learning framework called TGMM that simultaneously learns to solve multiple GMM tasks using a shared transformer backbone. The learned models are empirically demonstrated to effectively mitigate the limitations of classical methods such as Expectation-Maximization (EM) or spectral algorithms, at the same time exhibit reasonable robustness to distribution shifts. Theoretically, we prove that transformers can approximate both the EM algorithm and a core component of spectral methods (cubic tensor power iterations). These results bridge the gap between practical success and theoretical understanding, positioning transformers as versatile tools for unsupervised learning.  [^1] Code available at https://github.com/Rorschach1989/transformer-for-gmm
[…]
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various tasks in recent years. Transformers attention-is-all-you-need , the dominant architecture in modern LLMs BrownGPT3 , outperform many other neural network models in efficiency and scalability. Beyond language tasks, transformers have also demonstrated strong performance in other domains, such as computer vision survey-vision-TF ; TF-in-vision and reinforcement learning li2023TF-RL . Given their practical success, understanding the mechanisms behind transformers has attracted growing research interest. Existing studies often treat transformers as algorithmic toolboxes, investigating their ability to implement diverse algorithms (von-oswald23a, ; bai2023tfstats, ; lin2024transformers, ; giannou2025how, ; teh2025empiricalbayestransformers, )–a perspective linked to meta-learning hospedales2021meta .
@@
However, most research has focused on supervised learning settings, such as regression bai2023tfstats and classification giannou2025how , leaving the unsupervised learning paradigm relatively unexplored. Since transformer models are typically trained in a supervised manner, unsupervised learning poses inherent challenges for transformers due to the absence of labeled data. Moreover, given the abundance of unlabeled data in real-world scenarios, investigating the mechanisms of transformers in unsupervised learning holds significant implications for practical applications. The Gaussian mixture model (GMM) represents one of the most fundamental unsupervised learning tasks in statistics, with a rich historical background(GMM1969, ; Aitkin01081980, ) and ongoing research interest(gmm2021ACM-TOMM, ; dcgmm-nips-2021, ; loffler2021optimality, ; ndaoud2022sharp, ; gribonval2021statistical, ; yu2021scgmai, ). Two primary algorithmic approaches are existing for solving GMM problems: (1) likelihood-based methods employing the Expectation-Maximization (EM) algorithm(EM1977, ; BinYu-EM-2017, ), and (2) moment-based methods utilizing spectral algorithms(10.1145LearningGMM, ; JMLR:v15:anandkumar14b, ). However, both algorithms have inherent limitations. The EM algorithm is prone to convergence at local optima and is highly sensitive to initialization moitra2018algorithmic ; chijinEM . In contrast, while the spectral method is independent of initialization, it requires the number of components to be smaller than the data’s dimensionality—an assumption that restricts its applicability to problems involving many components in low-dimensional GMMs10.1145LearningGMM . In this work, we explore transformers for GMM parameter estimation. We empirically find that meta-trained transformers exhibit strong performance on GMM tasks without the aforementioned limitations. Notably, we construct transformer-based solvers that efficiently solve GMMs with varying component counts simultaneously. The experimental phenomena are further backed up by novel theoretical establishments: We prove that transformers can effectively learn GMMs with different components by approximating both the EM algorithm and a key component of spectral methods on GMM tasks.
Our contributions are mainly summarized as follows.
@@
  • We prove that ReLU-activated transformers can exactly approximate cubic tensor power iterations – a crucial component of spectral algorithms for GMM. This result stems from a novel insight about attention heads: they can perform computations along dimensions beyond the query-key-value framework. To the best of our knowledge, this is the first theoretical demonstration of transformers’ capacity for high-order tensor calculations.
  • Our work is among the earliest studies to investigate the mechanism of transformers in unsupervised learning settings. Our findings enhance the theoretical understanding of transformers and offer valuable insights for future research.
The paper is organized as follows. In Section 2, some background knowledge is introduced. In Section 3, we present the experimental details and findings. The theoretical results are proposed in Section 4, and some discussions are given in Section 5. The proofs and additional experimental results are given in the appendix.
@@
Related works. Recent research has explored the mechanisms by which transformers can implement various supervised learning algorithms. For instance, akyrek2023what , von-oswald23a , and bai2023tfstats demonstrate that transformers can perform gradient descent for linear regression problems in-context. lin2024transformers shows that transformers are capable of implementing Upper Confidence Bound (UCB) algorithms, as well as other classical algorithms in reinforcement learning tasks. giannou2025how reveals that transformers can execute in-context Newton’s method for logistic regression problems. teh2025empiricalbayestransformers illustrates that transformers can approximate Robbins’ estimator and solve Naive Bayes problems. kim2024transformers studies the minimax optimality of transformers on nonparametric regression.
@@
Comparison with prior theoretical works in unsupervised learning setting. Several recent studies have investigated the mechanisms of transformer-based models in mixture model settings he2025learningspectralmethodstransformers ; jin2025incontextlearningmixturelinear ; he2025transformersversusemalgorithm . Among these, he2025learningspectralmethodstransformers establishes that transformers can implement Principal Component Analysis (PCA) and leverages this to GMM clustering. However, their analysis is limited to the two-component case, restricting its broader applicability. The paper jin2025incontextlearningmixturelinear investigates the in-context learning capabilities of transformers for mixture linear models, a setting that differs from ours. Furthermore, their approximation construction of the transformer is limited to two-component GMMs, leaving the general case unaddressed. While they assume ReLU as the activation function–contrary to the conventional choice of softmax–their theoretical proofs rely on a key lemma from prior workpathak2024transformers that assumes softmax activation, thereby introducing an inconsistency in their assumptions. The paper he2025transformersversusemalgorithm studies the performance of transformers on multi-class GMM clustering, a setting closely related to ours. However, our work focuses on parameter estimation rather than clustering. From a technical standpoint, their approximation bound scales exponentially with the dimension $d$, and their analysis requires the number of attention heads $M$ to grow to infinity to achieve a meaningful bound. In contrast, our results are significantly stronger: our bounds depend polynomially on $d$, and we only require $M=O(1)$. From an empirical perspective, their experiments are conducted on a small-scale transformer, which fails to validate their theoretical claims.
@@
Notations. We introduce the following notations. Let $[n]\mathrel{\mathop{:}}={\{1,2,\cdots,n\}}$. All vectors are represented as column vectors unless otherwise specified. For a vector $v\in\mathbb{R}^{d}$, we denote $\left\|{v}\right\|$ as its Euclidean norm. For two sequences $a_{n}$ and $b_{n}$ indexed by $n$, we denote $a_{n}=O(b_{n})$ if there exists a universal constant $C$ such that $a_{n}\leq Cb_{n}$ for sufficiently large $n$.
@@ 2.1 Preliminaries
The Gaussian mixture model (GMM) is a cornerstone of unsupervised learning in statistics, with deep historical roots and enduring relevance in modern research. Since its early formalizations (GMM1969, ; Aitkin01081980, ), GMM has remained a fundamental tool for clustering and density estimation, widely applied across diverse domains. Recent advances have further explored the theoretical foundations of Gaussian Mixture Models (GMMs) (loffler2021optimality, ; ndaoud2022sharp, ; gribonval2021statistical, ), extended their applications in incomplete data settings (gmm2021ACM-TOMM, ), and integrated them with deep learning frameworks (dcgmm-nips-2021, ; yu2021scgmai, ). Due to their versatility and interpretability, GMMs remain indispensable in unsupervised learning, effectively bridging classical statistical principles with modern machine learning paradigms. We consider the (unit-variance) isotropic Gaussian Mixture Model with $K$ components, with its probability density function as
@@
where $\phi(x;\mu)$ is the standard Gaussian kernel, i.e. $\phi(x;\mu)=\frac{1}{(2\pi)^{d/2}}\exp\left(-\frac{1}{2}(x-\mu)^{\top}(x-\mu)\right)$. The parameter ${\boldsymbol{\theta}}$ is defined as ${\boldsymbol{\theta}}={\boldsymbol{\pi}}\cup{\boldsymbol{\mu}}$, where ${\boldsymbol{\pi}}\mathrel{\mathop{:}}={\{\pi_{1},\pi_{2},\cdots,\pi_{K}\}}$, $\pi_{k}\in\mathbb{R}$ and ${\boldsymbol{\mu}}={\{\mu_{1},\mu_{2},\cdots,\mu_{K}\}},\mu_{k}\in\mathbb{R}^{d}$, $k\in[K]$. We take $N$ samples ${\mathbf{X}}={\{X_{i}\}}_{i\in[N]}$ from model Eq. 1. ${\{X_{i}\}}_{i\in[N]}$ can be also rewritten as
@@
where ${\{y_{i}\}}_{i\in[N]}$ are iid discrete random variables with ${\mathbb{P}}{\left(y=k\right)}=\pi_{k}$ for $k\in[K]$ and ${\{Z_{i}\}}_{i\in[N]}$ are iid standard Gaussian random vector in $\mathbb{R}^{d}$.
@@
The EM algorithm EM1977 remains the most widely used approach for GMM parameter estimation. Due to space constraints, we propose the algorithm in Appendix A. Alternatively, the spectral algorithm (10.1145LearningGMM, ) offers an efficient moment-based approach that estimates parameters through low-order observable moments. A key component of this method is cubic tensor decomposition JMLR:v15:anandkumar14b . For brevity, we defer the algorithmic details to Appendix A.

### v2

@@ Abstract
The transformer architecture has demonstrated remarkable capabilities in modern artificial intelligence, among which the capability of implicitly learning an internal model during inference time is widely believed to play a key role in the understanding of pre-trained large language models. However, most recent works have been focusing on studying supervised learning topics such as in-context learning, leaving the field of unsupervised learning largely unexplored. This paper investigates the capabilities of transformers in solving Gaussian Mixture Models (GMMs), a fundamental unsupervised learning problem through the lens of statistical estimation. We propose a transformer-based learning framework called Transformer for Gaussian Mixture Models (TGMM) that simultaneously learns to solve multiple GMM tasks using a shared transformer backbone. The learned models are empirically demonstrated to effectively mitigate the limitations of classical methods such as Expectation-Maximization (EM) or spectral algorithms, at the same time exhibit reasonable robustness to distribution shifts. Theoretically, we prove that transformers can efficiently approximate both the Expectation-Maximization (EM) algorithm and a core component of spectral methods—namely, cubic tensor power iterations. These results not only improve upon prior work on approximating the EM algorithm, but also provide, to our knowledge, the first theoretical guarantee that transformers can approximate high-order tensor operations. Our study bridges the gap between practical success and theoretical understanding, positioning transformers as versatile tools for unsupervised learning.  [^1] Code available at https://github.com/Rorschach1989/transformer-for-gmm
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various tasks in recent years. Transformers(Vaswani et al., 2017), the dominant architecture in modern LLMs(Brown et al., 2020), outperform many other neural network models in efficiency and scalability. Beyond language tasks, transformers have also demonstrated strong performance in other domains, such as computer vision(Han et al., 2023; Khan et al., 2022) and reinforcement learning(Li et al., 2023a). Given their practical success, understanding the mechanisms behind transformers has attracted growing research interest. Existing studies often treat transformers as algorithmic toolboxes, investigating their ability to implement diverse algorithms(Von Oswald et al., 2023; Bai et al., 2023; Lin et al., 2024; Giannou et al., 2025; Teh et al., 2025)–a perspective linked to meta-learning(Hospedales et al., 2021).
@@
However, most research has focused on supervised learning settings, such as regression(Bai et al., 2023) and classification(Giannou et al., 2025), leaving the unsupervised learning paradigm relatively unexplored. Since transformer models are typically trained in a supervised manner, unsupervised learning poses inherent challenges for transformers due to the absence of labeled data. Moreover, given the abundance of unlabeled data in real-world scenarios, investigating the mechanisms of transformers in unsupervised learning holds significant implications for practical applications. The Gaussian mixture model (GMM) represents one of the most fundamental unsupervised learning tasks in statistics, with a rich historical background(DAY, 1969; Aitkin and Wilson, 1980) and ongoing research interest(Zhang et al., 2021; Manduchi et al., 2021; Löffler et al., 2021; Ndaoud, 2022; Gribonval et al., 2021; Yu et al., 2021). Two primary algorithmic approaches are existing for solving GMM problems: (1) likelihood-based methods employing the Expectation-Maximization (EM) algorithm(Dempster et al., 1977; Balakrishnan et al., 2017), and (2) moment-based methods utilizing spectral algorithms(Hsu and Kakade, 2013; Anandkumar et al., 2014). However, both algorithms have inherent limitations. The EM algorithm is prone to convergence at local optima and is highly sensitive to initialization(Moitra, 2018; Jin et al., 2016). In contrast, while the spectral method is independent of initialization, it requires the number of components to be smaller than the data’s dimensionality—an assumption that restricts its applicability to problems involving many components in low-dimensional GMMs(Hsu and Kakade, 2013).
In this work, we explore transformers for GMM parameter estimation to address two questions. (i) Can Transformers provably work for GMM in-context? (ii) Can Transformers empirically overcome the drawbacks of both EM algorithm and the spectral method? Our answers are affirmative. We find that meta-trained transformers exhibit strong performance on GMM tasks without the aforementioned limitations. Notably, we construct transformer-based solvers that efficiently solve GMMs with varying component counts simultaneously. The experimental phenomena are further backed up by novel theoretical establishments: We prove that transformers can effectively learn GMMs with different components by approximating both the EM algorithm and a key component of spectral methods on GMM tasks.
Main Contributions.
@@
  • We proved that transformers (with RELU activation) can implement cubic tensor power iterations- a crucial component of spectral algorithms for GMM. The proof is highly dependent on the multi-head structure of transformers. To the best of our knowledge, this is the first theoretical demonstration of transformers’ capacity for high-order tensor calculations.
@@
Related works. Recent research has explored the mechanisms by which transformers can implement various supervised learning algorithms. For instance, Akyürek et al. (2023), Von Oswald et al. (2023), and Bai et al. (2023) demonstrate that transformers can perform gradient descent for linear regression problems in-context. Lin et al. (2024) shows that transformers are capable of implementing Upper Confidence Bound (UCB) algorithms, as well as other classical algorithms in reinforcement learning tasks. Giannou et al. (2025) reveals that transformers can execute in-context Newton’s method for logistic regression problems. Teh et al. (2025) illustrates that transformers can approximate Robbins’ estimator and solve Naive Bayes problems. Kim et al. (2024) studies the minimax optimality of transformers on nonparametric regression. Some literature on density estimation using LLMs is discussed in Appendix˜A.
Comparison with prior theoretical works in unsupervised learning setting. Several recent studies have investigated the mechanisms of transformer-based models in mixture model settings(He et al., 2025a; Jin et al., 2024; He et al., 2025b). Among these, He et al. (2025a) establishes that transformers can implement Principal Component Analysis (PCA) and leverages this to GMM clustering. However, their analysis is limited to the two-component case, restricting its broader applicability.
@@
The paper Jin et al. (2024) investigates the in-context learning capabilities of transformers for mixture linear models, a setting that differs from ours. Furthermore, their approximation construction of the transformer is limited to two-component GMMs, leaving the general case unaddressed. While they assume ReLU as the activation function–contrary to the conventional choice of softmax–their theoretical proofs rely on a key lemma from prior work Pathak et al. (2024) that assumes softmax activation, thereby introducing an inconsistency in their assumptions. The paper He et al. (2025b) studies the performance of transformers on multi-class GMM clustering, a setting closely related to ours. However, our work focuses on parameter estimation rather than clustering. We give a discussion of our theoretical improvements over their work in detail in the following paragraph. From an empirical perspective, their experiments are conducted on a small-scale transformer, which fails to validate their theoretical claims.
Sharpness of our results. Our theoretical analysis fully leverages key architectural components of Transformers: the query-key-value mechanism, multi-head attention, and the properties of the activation function. It is worth pointing out that our result improves the prior work for EM approximation in several points: First, Our analysis shows that Transformers can approximate L-step EM algorithms with just O(L) layers, a significant improvement over prior work (He et al., 2025b) , which requires O(KL) layers (dependent on the number of components K). Second, unlike He et al. (2025b), which needs number of attention heads $M\rightarrow+\infty$ to get valid bounds, our results hold with $M=O(1)$, aligning better with real-world designs. Third, our approximation bounds scale polynomially in dimension $d$, unlike He et al. (2025b)’s exponential dependence–a crucial improvement for high-dimensional settings. We believe our results and proofs can offer profound insights for subsequent theoretical research on transformers.
Organization. The rest of paper is organized as follows. In Section˜2, some background knowledge is introduced. In Section˜3, we present the experimental details and findings. The theoretical results are proposed in Section˜4, and some discussions are given in Section˜5. The proofs and additional experimental results are given in the appendix.
@@
Notations. We introduce the following notations. Let $n]\mathrel{\mathop{:}}={\{1,2,\cdots,n\}}$. All vectors are represented as column vectors unless otherwise specified. For a vector $v\in\mathbb{R}^{d}$, we denote $\left\|{v}\right\|$ as its Euclidean norm. For two sequences $a_{n}$ and $b_{n}$ indexed by $n$, we denote $a_{n}=O(b_{n})$ if there exists a universal constant $C$ such that $a_{n}\leq Cb_{n}$ for sufficiently large $n$.
@@ 2.1 Preliminaries
The Gaussian mixture model (GMM) is a cornerstone of unsupervised learning in statistics, with deep historical roots and enduring relevance in modern research. Since its early formalizations(DAY, 1969; Aitkin and Wilson, 1980), GMM has remained a fundamental tool for clustering and density estimation, widely applied across diverse domains. Recent advances have further explored the theoretical foundations of Gaussian Mixture Models (GMMs)(Löffler et al., 2021; Ndaoud, 2022; Gribonval et al., 2021), extended their applications in incomplete data settings(Zhang et al., 2021), and integrated them with deep learning frameworks(Manduchi et al., 2021; Yu et al., 2021). Due to their versatility and interpretability, GMMs remain indispensable in unsupervised learning, effectively bridging classical statistical principles with modern machine learning paradigms. We consider the (unit-variance) isotropic Gaussian Mixture Model with $K$ components, with its probability density function as
@@
where $\phi(x;\mu)$ is the standard Gaussian kernel, i.e. $\phi(x;\mu)=\frac{1}{(2\pi)^{d/2}}\exp\left(-\frac{1}{2}(x-\mu)^{\top}(x-\mu)\right)$. The parameter $\boldsymbol{\theta}$} is defined as ${\boldsymbol{\theta}}={\boldsymbol{\pi}}\cup{\boldsymbol{\mu}}$, where ${\boldsymbol{\pi}}\mathrel{\mathop{:}}={\{\pi_{1},\pi_{2},\cdots,\pi_{K}\}}$, $\pi_{k}\in\mathbb{R}$ and ${\boldsymbol{\mu}}={\{\mu_{1},\mu_{2},\cdots,\mu_{K}\}},\mu_{k}\in\mathbb{R}^{d}$, $k\in[K]$. We take $N$ samples ${\mathbf{X}}={\{X_{i}\}}_{i\in[N]}$ from model Eq.˜1. ${\{X_{i}\}}_{i\in[N]}$ can be also rewritten as
@@
where ${\{y_{i}\}}_{i\in[N]}$ are i.i.d. discrete random variables with ${\mathbb{P}}{\left(y=k\right)}=\pi_{k}$ for $k\in[K]$ and ${\{Z_{i}\}}_{i\in[N]}$ are i.i.d. standard Gaussian random vector in $\mathbb{R}^{d}$.
@@
The EM algorithm(Dempster et al., 1977) remains the most widely used approach for GMM parameter estimation. Due to space constraints, we propose the algorithm in Appendix˜B. Alternatively, the spectral algorithm(Hsu and Kakade, 2013) offers an efficient moment-based approach that estimates parameters through low-order observable moments. A key component of this method is cubic tensor decomposition(Anandkumar et al., 2014). For brevity, we defer the algorithmic details to Appendix˜B.